Find the monthly mortgage payment for a mortgage of $93,000 at 7.75% per annum, amortized over 20 years.
Answer by ewatrrr(24785)


Hi
M = P [ i(1 + i)^n ] / [ (1 + i)^n - 1]
i = .0775/12 = .006458 (1 + i)n = (1.006458)^240 = 4.6877
M = 93000*.006458*4.6877/3.6877
Many mortgage calulators on line to check Your answer